The power supplied to a typical black-and-white television set is 88 W when the set is connected to 120 V.

(a) How much electrical energy does this set consume in 1 hour?
J

(b) A color television set draws about 2.7 A when connected to 120 V. How much time is required for it to consume the same energy as the black-and-white model consumes in 1 hour?
min
Answer #1

apply Power = eenrgy/time

sp energy E = Pt

E = 88 * 60 *60

E = 3.168 *10^5 Joules

------------------------

Power P = Vi = 2.7 *120

P = 324 Watts

so use Power = energy/time

so time here t = E/P

t = 3.168 e5/324

t = 977.77 secs

t = 16.29 mins
Answer #2

a)

Energy = Power * time

Energy = 88 * 3600

Energy = 316800 J

b) here ,

Energy = 2.7 * 120

Energy = 324 J

therefore,

324 * t =  316800

t = 977.8 s

t = 16.296 min

16.296 min time required for it to consume the same energy
Answer #3

a) P = 88 W = 88 J / s

Energy in 1 hr = 3600 sec = 88 * 3600 = 316800 J = 316.8 kJ

b) P = VI

= 120 * 2.7 = 324 W

E = Pt

E = 316800 J

316800 = 324 * t

t = 977.77 sec

= 16 minutes 18 seconds